# BOP-Distrib: Revisiting 6D Pose Estimation Benchmarks for Better Evaluation under Visual Ambiguities

### v2

@@ Results.
Table 1 reports the results. We can observe a large impact of the ground truth on the methods results (see Loss column) and ranking. In terms of ranking, the well established pose estimation task face the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods. For pose estimation of unseen objects, the difference of ranking can reach 6 places.
We observe large impacts on the recall performances as well. Whereas 29 of the 30 contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 6 methods with the original annotations, only 4 methods exceed this score with our annotations, with a majority of methods having less than 65%. Similarly, most of the contenders for the pose estimation of unseen objects task have a mean recall that exceeds 50% with the original annotations whereas only the 2 best RGB-D methods exceed this score once evaluated with our annotations. As shown in Figure 1 (Case 3), the image-wise annotation rejects poses when the image is not ambiguous. The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary material for visualizations).
@@ 5.4 Pose Distribution Evaluation
Experimental protocol
Similarly to Section 5.3, the methods were trained on the train set of T-LESS. The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discrete set of 6D pose obtained with the 6D pose annotation process introduced in Section 3. The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4. More details on methods processing and results formatting are given in the supplementary material.
Results
@@ 6 Limitations
While our solution improves the pose annotations, the resulting pose distribution depends on some parameters (tolerance factors, sampling of the SE(3) space, etc.). We selected their values carefully. We invite the reader to check the visualizations in the supplementary material.
@@ 7 Conclusion
1 Analysis of our BOP-D Annotations
1.1 Additionnal BOP-D Ground Truths Visualizations
@@
These images are taken from a video compilation of all ground truth annotations sorted by object identifier, also provided as supplementary material (BOP_D_id239_supp_newGT_visualizations.mp4), to convince the reader of their quality. We invite the reader to stop on some frames and check that the distribution recovered by our method does correspond to the ambiguities in the image for the object in the bounding box.
1.2 Differences between Original Annotations and Ours
@@
Figure 4 visualizes the differences between the poses that are accepted by BOP and the ones accepted when using our annotations. For this purpose, for each object, we provide a bar plot. This bar plot shows, for each image where the object appears, the percentage of poses considered correct by BOP that are also considered correct with our annotations. The bar plots show the percentages after sorting them, i.e., the bar on the left corresponds to the image with the smallest difference.
@@
The annotations for some objects remain mostly unchanged. But because our method analyses more finely the possible object symmetries, many poses are actually not accepted with our annotations for the other objects.
@@
Figure 4: Visualization of annotation changes compared to the T-LESS original annotations. Top: T-LESS objects with their identifier. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). T-LESS assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations. For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects. Our annotations significantly change the original annotations. The objects annotated as ’circular’ in BOP are highlighted in orange.
@@
Figure 5: Visualizing some BOP-D ground truths as computed by our method. Each example features an object of interest, in the bounding box in the left image, and its BOP-D pose distribution, split between the rotation part (center) and translation part (right). As no object present symmetries in translation here, the translation part of the distribution is always on the same point of $\mathbb{R}^{3}$. We provide much more examples in the accompanying video.
@@ 2.1 T-LESS Pose Estimation and Unseen Objects Pose Estimation
Recently, the BOP challenge [42] made public the pose estimations results of the methods evaluated in the leaderboard  [^3] https://bop.felk.cvut.cz/leaderboards/. Based on these results and the BOP toolkit  [^4] https://github.com/thodan/bop_toolkit/tree/master in which we implemented the variations of MSSD and MSPD, that use our per-image symmetries patterns instead of BOP global object symmetries, we were able to reprocess these pose estimates against our new ground truths. Our results have been presented in Table 1. Section 3 of supplementary material illustrates some of the failure cases with the new and more accurate ground truth.
@@ 2.2 Computation of the SpyroPose [9] Distribution Results
The implementation provided by the authors of SpyroPose allows to train a network for one object of T-LESS. We batched the training stage for all objects, and then batched the inference. For one object, SpyroPose produces more than 100 000 estimates, sorted by their probabilities, as it samples SO(3) $\times$ $\mathbb{R}^{3}$ (using a slice of $\mathbb{R}^{3}$). As the probabilities of these estimates quickly tend to zero, we reduce their distributions to the 400 best estimates. These 400 estimates are used for the evaluations of Section 5.4.
As no pose distribution evaluation existed, the authors of [9] reported only per-object averaged log-likelihood against BOP original ground truth.
@@ 2.3 Computation of the Lie-Pose Diffusion [19] Distribution Results
The implementation provided by the authors of Lie-Pose allows to train a network for all objects of T-LESS. The inference phase produces one pose estimate per image crop. We batched the inference phase, with varying seeds. We then merged all these estimates into a single set of poses per image crop. The authors report 1000 runs to produce a distribution. For our evaluation, we ran the code 100 times with different noises to produce Lie-Pose distribution results.
As no pose distribution evaluation existed, the authors of [19] reported only few illustrations of pose distributions results on T-LESS. They evaluated their method as a single pose estimation method, with a single run of the method.
@@ 3 Illustrations of Single Pose Re-evaluation
To do so, we look at the pose estimates of the method GDRNPP v1.4. When we evaluated GDRNPP v1.4 estimates against our new ground truths, it changed the method ranking from rank 6 to rank 10 in Table 1. Even more interestingly, the metrics went down from an MSSD of 88.4 and an MSPD of 90.9 to an MSSD of 60.6 and an MSPD of 63.7.
Figure 6 illustrates why MSSD and MSPD changed for this method. It appears that, although GDRNPP v1.4 estimates were close to the ground truth pose, its rotations were not precise. When they are evaluated against a symmetries pattern that is not precise, the evaluation appears correct. Our new ground truth shows that GDRNPP v1.4 tends not to align correctly some of the objects. Hence the drop in performances.
@@
(a) BOP continuous symmetry pattern of Object 2 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 4.46. (b) Our visual symmetry pattern (the much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 21.62.
(c) Corresponding image for (a) and (b): T-LESS Scene 1, Image 17, Object 2 (in bounding box).
@@
(d) BOP continuous symmetry pattern for Object 1 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 9.43. (e) Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 32.87.
@@
BOP continuous symmetry pattern of Object 4 (envelop) and GDRNPP v1.4 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 3.33. Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 35.38.
(h) Corresponding image for T-LESS Scene 5, Image 70, Objects 1 (d-e) and and 4 (f-g), in bounding boxes).
Figure 6: Impact of our annotations on Single Pose evaluation. We show here cases where the estimates by a state-of-the-art method (GDRNPP v1.4) produces fairly good MSSD errors when considering ground truth provided by BOP. For these cases, our more accurate ground truth yields worse MSSD errors, as it appears that the estimate belongs to the global symmetry pattern, but does not explain what is visible in the image.
@@ 4 Visualizing results by SpyroPose [9]
We display some of SpyroPose distribution estimates against our ground truth in Figure 7. For the case of the three instances of Object 1, SpyroPose correctly retrieves the single mode for Instance 1. The continuous symmetry of Instance 2 is partially retrieved.
Figure 7: Illustration of SpyroPose [9] results.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.

### v3

@@ Results.
Table 1 reports the results. We can observe a large impact of the ground truth on the methods results (see Loss column) and ranking. In terms of ranking, the well established pose estimation task faces the biggest changes, with gdrnppdet-pbrrealmegapose-multihyp moving from the 12th to the 3rd position for RGB methods and surfemb-pbr-rgbd-lin [8] moving from the 16th to the 5th place for the RGB-D methods. For pose estimation of unseen objects, the difference of ranking can reach 6 places.
We observe large impacts on the recall performances as well. Whereas 29 of the 30 contenders of pose estimation task have a mean recall up to 80% and even up to 90% for 6 methods with the original annotations, only 4 methods exceed this score with our annotations, with a majority of methods having less than 65%. Similarly, most of the contenders for the pose estimation of unseen objects task have a mean recall that exceeds 50% with the original annotations whereas only the 2 best RGB-D methods exceed this score once evaluated with our annotations. As shown in Figure 1 (Case 3), the image-wise annotation rejects poses when the image is not ambiguous. The Loss is explained by inaccurate pose estimates that were mistakenly validated by previous object-wise annotations (see supplementary materials for visualizations).
@@ 5.4 Pose Distribution Evaluation
Experimental protocol.
Similarly to Section 5.3, the methods were trained on the train set of T-LESS. The ground truth used for this dataset is the same than the one of Section 5.3 and is constituted of a discrete set of 6D pose obtained with the 6D pose annotation process introduced in Section 3. The performances of the methods are evaluated in Precision and Recall, using accuracy measures MPD and MSD as introduced in Section 4. More details on methods and results processing are given as supplementary material.
Results.
@@ 7 Conclusion
1 Pseudo-code of Per-image Ground Truth Pose Annotation
In this section, we give the key elements of our annotation method with the following pseudo-codes.
1# Input: CAD model, candidate transformation set, threshold
2# Output: per point epsilon sym set
3def EpsSym(ModelCAD, TransformSet, $\epsilon$, $\zeta$, resolution):
4 SampledModel = uniform_resampling(ModelCAD, resolution)
5 for point in SampledModel:
6 for T in TransformSet:
7 # Texture-less case
8 if testColor == False:
9 if knnSearch(T*point, SampledModel, $\epsilon$) == True:
10 EpsSym[point].append(T)
11 # Textured case
12 else:
13 if knnSearch(T*point, SampledModel, $\epsilon$) == True:
14 if $d_{Color}$(point, T*point) < $\zeta$:
15 EpsSym[point].append(T)
16
17 return EpsSym, SampledModel
[…]
The geometric ($d_{Geom}$) and colorimetric ($d_{Color}$) distances are implemented as follows:
$d_{Geom}(x,m)=||x-m||_{2}$ and
$d_{Color}(x,m)<\zeta\Leftrightarrow\min\Big{(}|h(x)-h(m)|,\\ |h(x)-h(m)-360|,|h(x)-h(m)+360|\Big{)}<\zeta_{h},\\ \text{AND~{}}|s(x)-s(m)|<\zeta_{s},\\ \text{AND~{}}|v(x)-v(m)|<\zeta_{v}.$
with $h(.),s(.),v(.)$ being hue, saturation and value of the point, and $\zeta_{h},\zeta_{s},\zeta_{v}$, the respective hue, saturation and value thresholds.
1# Input: SampledModel, EpsSym, PoseGT, MaskVisib, $\tau$
2# Output: EpsSymImage
3def SoftIntersection(SampledModel, EpsSym, PoseGT, MaskVisib, $\tau$):
4 # Count the vertices that vote for a transform
5 for point in SampledModel:
6 if K*[$R_{gt}$, $T_{gt}$]*point in mask_visib:
7 for T in EpsSym[point]:
8 H[T]++
9 Sort(H)
10 # Count the vertices that vote for a transform
11 for i in size(H):
12 if H[0]-H[i] < $\tau$:
13 EpsSymImage.append(i)
14
15 return EpsSymImage
Listing 2: Image annotation, pseudo-code of equation (4) from section 3.2.
1# Input: ModelCAD, EpsSymImage, PoseGT, SensorDepth, $\delta$
2# Output: EpsSymImageGlobal
3def PostProcessAnnotateImage(ModelCAD, EpsSymImage, PoseGT, SensorDepth, $\delta$, threshold):
4 depthGT = render(ModelCAD, PoseGT)
5 distGT = depthToDistImage(depthGT)
6 visibleMaskGT = generateMask(depthGT, SensorDepth)
7 for Ti in EpsSymImage:
8 depthEst = render(ModelCAD, Ti)
9 distEst = depthToDistImage(depthEst)
10 visibleMaskEst = generateMask(depthEst, SensorDepth)
11 maskIntersection = intersection(visibleMaskGT, visibleMaskEst)
12 maskUnion = union(visibleMaskGT, visibleMaskEst)
13 nbPixelsOutlier += maskUnion-maskIntersection
14 nbPixelsOutlier += abs(distGt - distEst)[maskIntersection] >= $\delta$
15 if nbPixelsOutlier < threshold:
16 EpsSymImageGlobal.append(Ti)
17
18 return EpsSymImageGlobal
Listing 3: Image annotation depth post-processing, pseudo-code of equation (4) from section 3.3, based on VSD implementation from BOP toolkit.
Section 3.3 does not use the VSD metric, as VSD is normalized by the number of pixels in the union of the visible masks. In our case, we need an absolute score and not a relative one, so that the counting of outliers has a metric meaning and can represent the size of a minimal disambiguating element.
2 Analysis of our BOP-Distrib Annotations
2.1 Accuracy of the annotations
Figure 4 illustrate renderings of all poses of a symmetries pattern.
Figure 4: Variation between poses in a single mode symmetries pattern. We display the renderings of the symmetries pattern of object 2 in image 7, scene 1. Although the SO(3) visualization may show spread points, these rotations result in similar poses.
2.2 False visible points and false occluded points detected by the pruning stage
[…]
2.3 Additionnal BOP-Distrib Ground Truths Visualizations
@@
These images are taken from a video compilation of all ground truth annotations sorted by object identifier, also provided as supplementary material (BOP_Distrib_id8513_supp_newGT_visualizations.mp4), to convince the reader of their quality. We invite the reader to stop on some frames and check that the distribution recovered by our method does correspond to the ambiguities in the image for the object in the bounding box.
2.4 T-LESS [13] Annotation Details
[…]
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 2.5$\times$2.5mm2, $\delta=5mm$ and $\tau_{pix}=30pix$.
2.5 YCB-V [48] Annotation Details
[…]
The object surface visibility ${\cal V}(M)$ is computed by Z-buffering using ground truth pose $P_{\text{GT}}$ and the 3D model of the object. For the robust symmetry pattern intersection, the soft intersection tolerance factor $\tau$ was experimentally adjusted to 28 3D points, resulting in a minimal disambiguating element size of roughly 5.3$\times$5.3mm2. The pruning stage was not necessary for YCB-V, as objects are simpler, with much less occlusions.
2.6 Differences between Original Object-wise BOP Annotations and Our Image-wise Annotations
@@
Figure 7 for T-LESS [13] and Figure 8 for YCB-V [48] highlight the differences between the poses that are accepted by BOP and the ones accepted when using our annotations. For this purpose, for each object, we provide a bar plot. This bar plot shows, for each image where the object appears, the percentage of poses considered correct by BOP that are also considered correct with our annotations. The bar plots show the percentages after sorting them, i.e., the bar on the left corresponds to the image with the smallest difference.
@@
Concerning T-LESS, the annotations for some objects remain mostly unchanged. But because our method analyses more finely the possible object symmetries, many poses are actually not accepted with our annotations for the other objects (mainly the ’circular’ ones). T-LESS is a very good dataset for our per-image pose distribution annotation method as it features objects with complex symmetries as well as a lot inter-object occlusions.
Concerning YCB-V, some objects such as the mug (object 14 on Figure 8) could yield very interesting symmetries with occlusions [30]. However, the disambiguating handle of the mug is never occluded. Hence the BOP symmetries that tag the mug as unambiguous. Overall, YCB-V has less potential for displaying visual ambiguities. Most objects are disambiguated by texture and the few ambiguous YCB-V objects do not face sufficient occlusions to become ambiguous (objects 11, 14 and 18 mainly), as depicted by the visualization of the scenes in Figure 9. We show here that our per-image annotation method is able to retrieve finer symmetries patterns, which have an effect when evaluating Single Pose Estimation methods as in Table 1.
@@
Figure 7: Visualization of annotation changes compared to the T-LESS original annotations. Top: T-LESS objects with their identifiers. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). T-LESS assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 18 is perfectly symmetrical and our method retrieves the same poses as the original annotations. For the other objects, in particular the objects with complex symmetry patterns like the first 4 objects, our annotations significantly change the original annotations. These changes in the annotations explain the score changes for T-LESS in Table 1. The objects annotated as ’circular’ in BOP are highlighted in orange.
Figure 8: Visualization of annotation changes compared to the YCB-V original annotations. Top: YCB-V objects with their identifiers. Bottom: For each object, we plot the percentages of poses kept from the original annotations by our method over the images (sorted by percentages). YCB-V assumes full rotational symmetries, while our annotation method captures more complex symmetry patterns. Only Object 13 is perfectly symmetrical, both in terms of geometry and texture, and our method retrieves the same poses as the original annotations. Object 18 is given as completely symmetrical but our method tags it as non-ambiguous due to its texture. Similarly, objects 1, 16, 19, 20 and 21 have few symmetrical poses for BOP annotations where our method keep always only one pose, as the texture disambiguate them. These changes in the annotations explain the score changes for YCB-V in Table 1. The objects annotated as ’circular’ in BOP are highlighted in orange.
Figure 9: Visualization of YCB-V scenes. The scenes present at most 6 different objects, with one instance of each. The objects with potential ambiguities (mainly objects 14 and 18) have limited occlusions. YCB-V is not really suited for pose distribution evaluation as no object becomes ambiguous, apart from object 13.
Figure 10: Visualization of T-LESS scenes. The scenes present up to 18 instances of different objects, with a lot of occlusions and with ambiguous objects. T-LESS is really suited for pose distribution evaluation. (illustration borrowed from [13])
@@
Figure 11: Visualizing some BOP-Distrib ground truths as computed by our method. Each example features an object of interest, in the bounding box in the left image, and its BOP-Distrib pose distribution, split between the rotation part (center) and translation part (right). As no object present symmetries in translation here, the translation part of the distribution is always on the same point of $\mathbb{R}^{3}$. We provide much more examples in the accompanying video.
@@ 3.1 T-LESS Pose Estimation and Unseen Objects Pose Estimation
Recently, the BOP challenge [43] made public the pose estimation results of the methods evaluated in the leaderboard  [^3] https://bop.felk.cvut.cz/leaderboards/. Based on these results and the BOP toolkit  [^4] https://github.com/thodan/bop_toolkit/tree/master in which we implemented the variations of MSSD and MSPD, that use our per-image symmetries patterns instead of BOP global object symmetries, we were able to reprocess these pose estimates against our new ground truths. Our results on T-LESS have been presented in Table 1. Section 4 of supplementary material illustrates some of the failure cases with the new and more accurate ground truth.
3.2 YCB-V Pose Estimation
We conducted a similar experience of reprocessing BOP competitors on our re-annotation of the YCB-V. Our results have been presented in Table 1.
@@ 3.3 Computation of the SpyroPose [9] Distribution Results
Beyond finer evaluation of Single Pose Estimation methods, our new per-image annotations allow us to propose the first evaluation on real data of Pose Distribution Estimation methods. As no pose distribution evaluation existed, the authors of [9] reported only per-object averaged log-likelihood against BOP original ground truth.
The implementation provided by the authors of SpyroPose allows to train a network for one object of T-LESS. We batched the training stage for all objects, and then batched the inference. For one object, SpyroPose produces more than 100 000 estimates, sorted by their probabilities, as it samples SO(3) $\times$ $\mathbb{R}^{3}$ (using a slice of $\mathbb{R}^{3}$). As the probabilities of these estimates quickly tend to zero, we reduce their distributions to the 400 best estimates. These 400 estimates are used for the evaluations of Section 5.4 of the article.
@@ 3.4 Computation of the Lie-Pose Diffusion [19] Distribution Results
Similarly, as no pose distribution evaluation existed, the authors of [19] reported only few qualitative illustrations of pose distribution results on T-LESS. They evaluated their method as a Single Pose Estimation method, with a single run of the method.
The implementation provided by the authors of Lie-Pose allows to train a network for all objects of T-LESS. The inference phase produces one pose estimate per image crop. We batched the inference phase, with varying seeds. We then merged all these estimates into a single set of poses per image crop. The authors report 1000 runs to produce a distribution. For our evaluation, due to important computation time, we ran the code 100 times with different input noises to produce Lie-Pose distribution results.
@@ 4 Illustrations of Single Pose Re-evaluation
To do so, we look at the pose estimates of the method gdrnpp-pbrreal-rgbd-mmodelv1.3. When we evaluated gdrnpp-pbrreal-rgbd-mmodelv1.3 estimates against our new ground truths, it changed the method ranking from rank 7 to rank 10 in Table 1. Even more interestingly, the metrics went down from an MSSD of 88.4 and an MSPD of 90.9 to an MSSD of 60.6 and an MSPD of 63.7.
Figure 12 illustrates why MSSD and MSPD changed for this method. It appears that, although gdrnpp-pbrreal-rgbd-mmodelv1.3 estimates were close to the ground truth poses, its rotations were not precise. When they are evaluated against a symmetries pattern that is not precise, the evaluation appears correct. Our new ground truth shows that gdrnpp-pbrreal-rgbd-mmodelv1.3 tends not to align correctly some of the objects. Hence the drop in performances.
@@
(a) BOP continuous symmetry pattern of Object 2 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 4.46. (b) Our visual symmetry pattern (the much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 21.62.
@@
(c) Corresponding image for (a) and (b): T-LESS Scene 1, Image 17, Object 2 (in bounding box) and rendering of the pose (red arrow highlights disambiguating element).
@@
(d) BOP continuous symmetry pattern for Object 1 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 9.43. (e) Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 32.87.
Refer to caption Refer to caption
BOP continuous symmetry pattern of Object 4 (envelop) and gdrnpp-pbrreal-rgbd-mmodelv1.3 estimate (plain circle). The circle belongs to the envelop, yielding a low MSSD error of 3.33. Our visual symmetry pattern (much smaller envelop) and GDRNPP estimate (plain circle). The circle does not belong to the envelop anymore, the MSSD error becomes 35.38.
Refer to caption Refer to caption Refer to caption
(h) Corresponding image for T-LESS Scene 5, Image 70, Objects 1 (d-e) and and 4 (f-g), in bounding boxes) and renderings of the poses (red arrow highlights disambiguating elements).
Figure 12: Impact of our annotations on Single Pose evaluation. We show here cases where the estimates by a state-of-the-art method (gdrnpp-pbrreal-rgbd-mmodelv1.3) produces fairly good MSSD errors when considering ground truth provided by BOP. For these cases, our more accurate ground truth yields worse MSSD errors, as it appears that the estimate belongs to the global symmetry pattern, but does not explain what is visible in the image.
@@ 5 Visualizing results by SpyroPose [9]
We display some of SpyroPose distribution estimates against our ground truth in Figure 13 on T-LESS. For the case of the three instances of Object 1, SpyroPose correctly retrieves the single mode for Instance 1. The continuous symmetry of Instance 2 is partially retrieved.
Figure 13: Illustration of SpyroPose [9] results on T-LESS.. We show the distribution estimates for the 3 instances of Object 1 (in the bounding box). The ground truth distribution is displayed as an envelop for the rotation part and as red star for the translation part.